Suppose that g(3)=1, g'(3)=2 and f'(-1)= 0.5 find f(g(x))'3


1
Expert's answer
2021-01-05T12:54:02-0500

I think the question is wrong. It will be either "g(3)=-1" or "f'(1)=0.5." Let "F(x)=f(g(x))" .

Then "F'(x)=\frac {d}{dx}[f(g(x))]"

"=f'(g(x)).\frac{d}{dx}[g(x)]"

"=f'(g(x)).g'(x)"

Therefore, "F'(3)=f'(g(x))(3)"

"=f'(g(3)).g'(3)"

"=2.f'(-1)" [ taking "g(3)=-1" ]

"=2.\frac{1}{2}" [As "f'(-1)=0.5" ]

"=1"

"\therefore f(g(x))'(3)=1."